OpenStudy (anonymous):

Solve the system of equations using Gaussian elimination or gauss-Jordan elimination.

OpenStudy (anonymous):

5x+4y+z=3 4x-3y-z=37 5x+y+2z=16

OpenStudy (anonymous):

Add equation 1 and 2 together, the z terms subtract out. Multiply the second equation by 2 and add the result to the third equation, the z terms will subtract out. You now have two equations in two unknowns.

OpenStudy (anonymous):

I need to do this matrix way not that way I know the answer but I need to learn how to do it matrix

OpenStudy (anonymous):

@Hero help please

OpenStudy (anonymous):

@ParthKohli help please

OpenStudy (anonymous):

I know the answers is x= 5, y=-5, z=-2

OpenStudy (anonymous):

I didn't understand how my teacher did it so I am ask for help here

OpenStudy (anonymous):

I need to do it matrix way till the end

OpenStudy (anonymous):

@twvogels